OpenStudy (anonymous):

A circle has its center at (-2, 5) and a radius of 4 units. What is the equation of the circle? answer choices: (x + 2)2 + (y - 5)2 = 16 (x + 2)2+ (y + 5)2= 16 (x + 2)2+(y - 5)2 = 4 (x - 2)2+(y + 5)2 = 4

OpenStudy (anonymous):

either a or b, i wanna say b

OpenStudy (anonymous):

no you don't

OpenStudy (anonymous):

is be supose to say (x-2)?

OpenStudy (anonymous):

suppose*

OpenStudy (anonymous):

a) was the right choice \[(x-h)^2+(y-k)^2=r^2\] with \(h=-2, k=5, r=4\)

OpenStudy (anonymous):

make the replacement and get \[(x-(-2))^2+(y-5)^2=4^2\] \[(x+2)^2+(y-5)^2=16\]

OpenStudy (shane_b):

You can just look at it and see that it can't be b or d based on the signs of the constants in the parentheses. Looking at the other side of the equation, you can eliminate c as well.

OpenStudy (anonymous):

Sorry its a! i meant to put a!